An investor buys $17 thousand dollars of ABT stock at $20 per share, using 50% initial margin. The broker charges 9% APR compounded daily on the loan, and requires a 35% maintenance margin. The stock pays $0.59 per share dividend each year. If the stock is sold at the end of the year at $22 per share, what is the investor's rate of return?
Answer #1

Investor's equity at the beginning = $17,000 * 50% = $8,500.

Total value at the end of the year = ($17,000 / 20) * ($22 + $0.59)
= 850 * $22.59
= $19,201.50

Loan repayment = 17,000 - 8,500 * (1 + 9%/365)^365
= 8,500 * 1.0942
= $9,300.38

Equity value at the end = Total value - loan repayment
= $19,201.50 - $9,300.38
= $9,901.12

Return of rate = ($9,901.12 / $8,500) - 1
= 1.1648 - 1
= 0.1648 or 16.48%

Investor's rate of return = 16.48%